A rocket is launched from a cliff and it can be represented by the following function. h(t)= -16t squared + 80t + 384, where x is the time in seconds and h is the height of the rocket. Identify the zeroes of the function.
What is the height of the rocket?

Answer by lwsshak3(11628)

A rocket is launched from a cliff and it can be represented by the following function. h(t)= -16t squared + 80t + 384, where x is the time in seconds and h is the height of the rocket. Identify the zeroes of the function.
What is the height of the rocket?
***
h(t)=-16t^2+80t+384
complete the square:
h(t)=-16(t^2-5t+25/4)+100+384
h(t)=-16(t-5/2)^2+484
This is an equation of a parabola that opens downward with vertex at (5/2, 384)
maximum height of the rocket is 384 units 2.5 sec after launching
..
-16t^2+80t+384=0
-t^2-5t-24=0
(-t-3)(t-8)=0
t=-3
or
t=8
zeros are: -3, 8
height of rocket depends on time after launch